# MAGPIE: Multilevel–Adaptive–Guided solver for ptychographic phase retrieval

### v4

@@ 6 Acknowledgments
$=\frac{1}{2}\sum_{k=1}^{N}\left[\left\|\bm{Q}P_{k}\bm{z}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)\right\|_{2}^{2}-\left\|\bm{Q}P_{k}\bm{z}^{(j)}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)\right\|_{2}^{2}\right]\,,$
$=\frac{1}{2}\sum_{k=1}^{N}\bigg{[}\left\|\bm{Q}P_{k}\bm{z}\right\|^{2}-\overline{\bm{Q}P_{k}\bm{z}}\cdot\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)-\bm{Q}P_{k}\bm{z}\cdot\overline{\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)}$
$\quad\quad\quad-\left\|\bm{Q}P_{k}\bm{z}^{(j)}\right\|^{2}+\overline{\bm{Q}P_{k}\bm{z}^{(j)}}\cdot\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)+\bm{Q}P_{k}\bm{z}^{(j)}\cdot\overline{\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)}\bigg{]}\,,$
$=\frac{1}{2}\sum_{k=1}^{N}\bigg{[}\left\|\bm{Q}P_{k}\bm{z}-\bm{Q}P_{k}\bm{z}^{(j)}\right\|^{2}+\overline{\bm{Q}P_{k}\bm{z}-\bm{Q}P_{k}\bm{z}^{(j)}}\cdot\left(\bm{Q}P_{k}\bm{z}^{(j)}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)\right)$
$\quad\quad\quad+\left(\bm{Q}P_{k}\bm{z}-\bm{Q}P_{k}\bm{z}^{(j)}\right)\cdot\overline{\bm{Q}P_{k}\bm{z}^{(j)}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)}\bigg{]}\,,$
@@
Appendix D Useful Lemmas
Lemma D.1.
@@ Appendix D Useful Lemmas
$\texttt{LHS}_{i}=\bm{A}_{i}(\bm{I}_{h}^{H}\bm{B})_{i}=\bm{A}_{i}\sum_{j}(\bm{I}_{h}^{H})_{ij}\bm{B}_{j}\,.$
$\texttt{RHS}_{i}$ $=(\bm{I}_{h}^{H}[(\bm{I}_{H}^{h}\bm{A})\odot\bm{B}])_{i}\,,$
$=\sum_{j}(\bm{I}_{h}^{H})_{ij}[(\bm{I}_{H}^{h}\bm{A})\odot\bm{B}]_{j}\,,$
$=\sum_{j}(\bm{I}_{h}^{H})_{ij}\left[\sum_{k}(\bm{I}_{H}^{h})_{jk}\bm{A}_{k}\right]\bm{B}_{j}\,,$
$=\sum_{k}\bm{A}_{k}\sum_{j}(\bm{I}_{h}^{H})_{ij}(\bm{I}_{H}^{h})_{jk}\bm{B}_{j}\,,$
$=\sum_{k}\bm{A}_{k}\sum_{j}(\bm{I}_{h}^{H})_{ij}4(\bm{I}_{h}^{H})_{kj}\bm{B}_{j}\,,$
$=\sum_{k}\bm{A}_{k}\sum_{j}(\bm{I}_{h}^{H})_{ij}\delta_{ik}\bm{B}_{j}\,,$
$=\bm{A}_{i}\sum_{j}(\bm{I}_{h}^{H})_{ij}\bm{B}_{j}\,.$
Lemma D.2.
For any $\bm{B}\in\mathbb{C}^{n_{H}^{2}}$ and any function $f:\mathbb{C}\to\mathbb{C}$, we define the operators
@@
to be the element-wise applications of $f$, i.e., for any input vector the operator applies $f$ to every entry. Then, for any $\bm{C}\in\mathbb{C}^{n_{H}^{2}}$,
$\bm{f}_{h}\left(\bm{I}_{H}^{h}\bm{C}\right)=\bm{I}_{H}^{h}\bm{f}_{H}(\bm{C})\,.$
Lemma D.3.
@@
Lemma D.2 and Lemma D.3 are straightforward to check by comparing corresponding entries of both sides, so we omit the proofs.
Lemma D.4.
@@
Appendix E Proof for Proposition 3.1
@@
We provide proofs to the statements in the same order.
  • By Lemma D.1 and Lemma D.2, we have
@@
    $=\bm{I}_{h}^{H}\left(\bm{I}_{H}^{h}\left(\frac{1}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)|\bm{Q}|^{2}\right)\,,$
    $=\left(\frac{1}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\bm{I}_{h}^{H}\left(|\bm{Q}|^{2}\right)\,,$
@@
    $\left|\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\right|=\frac{|(\bm{I}_{H}^{h}\bm{Q}_{H})||\bm{Q}|^{2}}{|\bm{Q}|\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}\right)}=\frac{\left(\bm{I}_{H}^{h}|\bm{Q}_{H}|\right)|\bm{Q}|}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\leq\frac{\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|\right)|\bm{Q}|}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,.$
    Then, by Lemma D.1, Lemma D.2, and Lemma D.3, we have
    $\bm{I}_{h}^{H}\left|\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\right|$ $\leq\bm{I}_{h}^{H}\left(\frac{\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|\right)|\bm{Q}|}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\,,$
    $=\bm{I}_{h}^{H}\left(\bm{I}_{H}^{h}\left(\frac{\bm{I}_{h}^{H}|\bm{Q}|}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)|\bm{Q}|\right)\,,$
    $=\left(\frac{\bm{I}_{h}^{H}|\bm{Q}|}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\bm{I}_{h}^{H}|\bm{Q}|\,,$
@@
    $\bm{I}_{h}^{H}\left|\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\right|\leq 1\implies\left|\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\right|\leq 4\,,$
    where the inequalities are satisfied element-wise. We note the sharp bound for the last inequality is $3/2$, which is much harder to find.
@@
Appendix F Proof for Proposition 3.2
@@
$\widetilde{\Phi}_{H,k}\left(\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}\right);\bm{z}_{k}^{(j)}\right)$ $=\frac{1}{2}\left\|\bm{Q}_{H}\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}^{2}\,,$
$=\frac{1}{2}\left\|\bm{I}_{h}^{H}\left((\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}^{2}\,,$
$=\frac{1}{2}\left\|\bm{I}_{h}^{H}\left(\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\bm{Q}\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}^{2}\,,$
$=\frac{1}{2}\left\|\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\bm{Q}\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}^{2}\,,$
$=\frac{1}{2}\left\|\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\left(\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}^{2}\,,$
$\leq\frac{1}{2}\left\|\bm{I}_{h}^{H}\right\|_{2}^{2}\left\|\bm{W}_{\mathcal{R}}\right\|_{\infty}^{2}\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}\,,$
@@
where the second equality follows from Lemma D.1, and the last inequality follows from Lemma D.4.
@@
$\left\|\nabla_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}\right);\bm{z}_{k}^{(j)}\right)\right\|_{2}$ $=\left\|\overline{\bm{Q}_{H}}\left(\bm{Q}_{H}\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}\,,$
$=\left\|\overline{\bm{Q}_{H}}\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\left(\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}\,,$
$=\left\|\bm{I}_{h}^{H}\left(\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\overline{\bm{Q}}\left(\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}\,,$
$=\left\|\bm{I}_{h}^{H}\left(\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}\,,$
$\leq\left\|\bm{I}_{h}^{H}\right\|_{2}\left\|\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\right\|_{\infty}\left\|\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)\right\|_{2}\,.$
Since $\left\|\bm{I}_{h}^{H}\right\|_{2}=\frac{1}{2}$ by Lemma D.4, it suffices to check $\left\|\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\right\|_{\infty}=\left\|\bm{W}_{\bm{u}}^{H}\right\|_{\infty}$ and it is bounded by 1. To this end, by Lemma D.2, Lemma D.3, we find
$\left\|\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\right\|_{\infty}$ $=\left\|\frac{\overline{\bm{I}_{H}^{h}\bm{Q}_{H}}}{\overline{\bm{Q}}}\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\right\|_{\infty}\,,$
$=\left\|\frac{|\bm{I}_{H}^{h}\bm{Q}_{H}|^{2}}{|\bm{Q}|^{2}}\frac{|\bm{Q}|^{2}}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right\|_{\infty}\,,$
@@
where the last equality is discussed in the proof of Lemma D.4. Sharpness of the second inequality is straightforward to check by setting both $\bm{Q}$ and $\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$ as vectors of ones, in which case $\nabla_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}\right);\bm{z}_{k}^{(j)}\right)$ is also a vector of ones. ∎
Appendix G Proof for Theorem 3.3
Proposition G.1.
@@
$\bm{e}_{H,k}=-\bm{A}\odot\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,,$
@@
Additionally, if $\bm{A}$ is positive, then equality implies $\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)=0$.
@@
$\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)^{*}\left(\bm{I}_{H}^{h}\bm{e}_{H,k}\right)$ $=-\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)^{*}\left(\bm{I}_{H}^{h}\bm{A}\odot\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\right)\,,$
$=-\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)^{*}\bm{I}_{H}^{h}\operatorname{diag}(\bm{A})\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,,$
$=-4\left(\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\right)^{*}\operatorname{diag}(\bm{A})\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,,$

### v6

@@ Appendix A Notational convention
$=\frac{1}{2}\sum_{k=1}^{N}\left[\left\|\bm{Q}\odot P_{k}\bm{z}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)\right\|_{2}^{2}-\left\|\bm{Q}\odot P_{k}\bm{z}^{(j)}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)\right\|_{2}^{2}\right]\,$,
$=\frac{1}{2}\sum_{k=1}^{N}\bigg[\left\|\bm{Q}\odot P_{k}\bm{z}\right\|^{2}-\overline{\bm{Q}\odot P_{k}\bm{z}}\cdot\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)-\bm{Q}\odot P_{k}\bm{z}\cdot\overline{\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)}$
$\quad\quad\quad-\left\|\bm{Q}\odot P_{k}\bm{z}^{(j)}\right\|^{2}+\overline{\bm{Q}\odot P_{k}\bm{z}^{(j)}}\cdot\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)+\bm{Q}\odot P_{k}\bm{z}^{(j)}\cdot\overline{\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)}\bigg]\,$,
$=\frac{1}{2}\sum_{k=1}^{N}\bigg[\left\|\bm{Q}\odot P_{k}\bm{z}-\bm{Q}\odot P_{k}\bm{z}^{(j)}\right\|^{2}+\overline{\bm{Q}\odot P_{k}\bm{z}-\bm{Q}\odot P_{k}\bm{z}^{(j)}}\cdot\left(\bm{Q}\odot P_{k}\bm{z}^{(j)}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)\right)$
$\quad\quad\quad+\left(\bm{Q}\odot P_{k}\bm{z}-\bm{Q}\odot P_{k}\bm{z}^{(j)}\right)\cdot\overline{\bm{Q}\odot P_{k}\bm{z}^{(j)}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)}\bigg]\,$,
@@
Appendix E Useful Lemmas
Lemma E.1.
@@ Appendix E Useful Lemmas
$\texttt{LHS}_{i}=\bm{B}_{i}(\bm{I}_{h}^{H}\bm{A})_{i}=\bm{B}_{i}\sum_{j}(\bm{I}_{h}^{H})_{ij}\bm{A}_{j}\,.$
$\texttt{RHS}_{i}$ $=(\bm{I}_{h}^{H}[(\bm{I}_{H}^{h}\bm{B})\odot\bm{A}])_{i}\,$,
$=\sum_{j}(\bm{I}_{h}^{H})_{ij}[(\bm{I}_{H}^{h}\bm{B})\odot\bm{A}]_{j}\,$,
$=\sum_{j}(\bm{I}_{h}^{H})_{ij}\left[\sum_{k}(\bm{I}_{H}^{h})_{jk}\bm{B}_{k}\right]\bm{A}_{j}\,$,
$=\sum_{k}\bm{B}_{k}\sum_{j}(\bm{I}_{h}^{H})_{ij}(\bm{I}_{H}^{h})_{jk}\bm{A}_{j}\,$,
$=\sum_{k}\bm{B}_{k}\sum_{j}(\bm{I}_{h}^{H})_{ij}4(\bm{I}_{h}^{H})_{kj}\bm{A}_{j}\,$,
$=\sum_{k}\bm{B}_{k}\sum_{j}(\bm{I}_{h}^{H})_{ij}\delta_{ik}\bm{A}_{j}\,$,
$=\bm{B}_{i}\sum_{j}(\bm{I}_{h}^{H})_{ij}\bm{A}_{j}\,.$
Lemma E.2.
For any function $f:\mathbb{C}\to\mathbb{C}$, we define the operators
@@
to be the element-wise applications of $f$, i.e., for any input vector, the operator applies $f$ to every entry. Then, for any $\bm{B}\in\mathbb{C}^{n_{H}^{2}}$,
$\bm{f}_{h}\left(\bm{I}_{H}^{h}\bm{B}\right)=\bm{I}_{H}^{h}\bm{f}_{H}(\bm{B})\,.$
Lemma E.3.
@@
Lemma E.2 and Lemma E.3 are straightforward to check by comparing corresponding entries of both sides, so we omit the proofs.
Lemma E.4.
@@
Appendix F Proof for Proposition 4.1
@@
We provide proofs for the statements in the same order.
  • By Lemma E.1 and Lemma E.2, we have
@@
    $=\bm{I}_{h}^{H}\left(\bm{I}_{H}^{h}\left(\frac{1}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\odot|\bm{Q}|^{2}\right)\,$,
    $=\left(\frac{1}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\odot\bm{I}_{h}^{H}|\bm{Q}|^{2}\,$,
@@
    $\left|\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}}{\bm{Q}}\right|=\frac{|(\bm{I}_{H}^{h}\bm{Q}_{H})|\odot|\bm{Q}|^{2}}{|\bm{Q}|\odot\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}\right)}=\frac{\left(\bm{I}_{H}^{h}|\bm{Q}_{H}|\right)\odot|\bm{Q}|}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\leq\frac{\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|\right)\odot|\bm{Q}|}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,.$
    Then, by Lemma E.1, Lemma E.2, and Lemma E.3, we have
    $\bm{I}_{h}^{H}\left|\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}}{\bm{Q}}\right|$ $\leq\bm{I}_{h}^{H}\left(\frac{\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|\right)\odot|\bm{Q}|}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\,$,
    $=\bm{I}_{h}^{H}\left(\bm{I}_{H}^{h}\left(\frac{\bm{I}_{h}^{H}|\bm{Q}|}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\odot|\bm{Q}|\right)\,$,
    $=\left(\frac{\bm{I}_{h}^{H}|\bm{Q}|}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\odot\bm{I}_{h}^{H}|\bm{Q}|\,$,
@@
    $\bm{I}_{h}^{H}\left|\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}}{\bm{Q}}\right|\leq 1\implies\left|\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}}{\bm{Q}}\right|\leq 4\,$,
    where the inequalities are satisfied element-wise. We note the sharp bound for the last inequality is $3/2$, but its more involved derivation is omitted here.
@@
Appendix G Proof for Proposition 4.2
@@
$\widetilde{\Phi}_{H,k}\left(\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\odot\bm{z}_{k}\right);\bm{z}_{k}^{(j)}\right)$ $=\frac{1}{2}\left\|\bm{Q}_{H}\odot\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\odot\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\odot\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}^{2}\,$,
$=\frac{1}{2}\left\|\bm{I}_{h}^{H}\left((\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}\odot\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\odot\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}^{2}\,$,
$=\frac{1}{2}\left\|\bm{I}_{h}^{H}\left(\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}}{\bm{Q}}\odot\bm{Q}\odot\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\odot\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}^{2}\,$,
$=\frac{1}{2}\left\|\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\odot\bm{Q}\odot\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\odot\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}^{2}\,$,
$=\frac{1}{2}\left\|\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\odot\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}^{2}\,$,
$\leq\frac{1}{2}\left\|\bm{I}_{h}^{H}\right\|_{2}^{2}\left\|\bm{W}_{\mathcal{R}}\right\|_{\infty}^{2}\left\|\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}\,$,
@@
where the second equality follows from Lemma E.1, and the last equality follows from Lemma E.4.
@@
$\left\|\nabla_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\odot\bm{z}_{k}\right);\bm{z}_{k}^{(j)}\right)\right\|_{2}$ $=\left\|\overline{\bm{Q}_{H}}\odot\left(\bm{Q}_{H}\odot\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\odot\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\odot\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}\,$,
$=\left\|\overline{\bm{Q}_{H}}\odot\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\odot\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}\,$,
$=\left\|\bm{I}_{h}^{H}\left(\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\odot\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\odot\overline{\bm{Q}}\odot\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}\,$,
$=\left\|\bm{I}_{h}^{H}\left(\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\odot\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\odot\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}\,$,
$\leq\left\|\bm{I}_{h}^{H}\right\|_{2}\left\|\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\odot\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\right\|_{\infty}\left\|\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)\right\|_{2}\,.$
Since $\left\|\bm{I}_{h}^{H}\right\|_{2}=\frac{1}{2}$ by Lemma E.4, it suffices to check $\left\|\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\odot\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\right\|_{\infty}=\left\|\bm{W}_{\bm{u}}^{H}\right\|_{\infty}$ and it is bounded by 1. To this end, by Lemma E.2, Lemma E.3, we find
$\left\|\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\odot\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\right\|_{\infty}$ $=\left\|\frac{\overline{\bm{I}_{H}^{h}\bm{Q}_{H}}}{\overline{\bm{Q}}}\odot\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}}{\bm{Q}}\right\|_{\infty}\,$,
$=\left\|\frac{|\bm{I}_{H}^{h}\bm{Q}_{H}|^{2}}{|\bm{Q}|^{2}}\odot\frac{|\bm{Q}|^{2}}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right\|_{\infty}\,$,
@@
where the last equality is discussed in the proof of Lemma E.4. The sharpness of the second inequality is straightforward to check by setting both $\bm{Q}$ and $\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$ as vectors of ones, in which case $\nabla_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\odot\bm{z}_{k}\right);\bm{z}_{k}^{(j)}\right)$ is also a vector of ones. ∎
Appendix H Proof for Theorem 4.3
Proposition H.1.
@@
$\bm{e}_{H,k}=-\bm{A}\odot\left(\bm{I}_{h}^{H}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,$,
@@
Additionally, if $\bm{A}$ is positive, then equality implies $\bm{I}_{h}^{H}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)=0$.
@@
$\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)^{*}\left(\bm{I}_{H}^{h}\bm{e}_{H,k}\right)$ $=-\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)^{*}\left(\bm{I}_{H}^{h}\bm{A}\odot\left(\bm{I}_{h}^{H}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\right)\,$,
$=-\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)^{*}\bm{I}_{H}^{h}\operatorname{diag}(\bm{A})\left(\bm{I}_{h}^{H}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,$,
$=-4\left(\left(\bm{I}_{h}^{H}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\right)^{*}\operatorname{diag}(\bm{A})\left(\bm{I}_{h}^{H}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,$,